OpenStudy (anonymous):

f(x) = 9 - 4x, g(x) = -2x + 4 Find (f + g)(x)

OpenStudy (akshay_budhkar):

(f+g)x=fx + gx so u just need to add those two functions you can do that?

OpenStudy (anonymous):

so 9-4x + -2x+4 , Im really confused with these problems

OpenStudy (akshay_budhkar):

yes

OpenStudy (akshay_budhkar):

add them

OpenStudy (akshay_budhkar):

what is 9+4?

OpenStudy (anonymous):

13

OpenStudy (akshay_budhkar):

ok now what is -2x + -4x ?

OpenStudy (anonymous):

-6x

OpenStudy (akshay_budhkar):

so your answer is -6x + 13 isn't that simple :D

OpenStudy (akshay_budhkar):

you get it or are you still confused?

OpenStudy (anonymous):

oh okay, yes i get it, There is a similar problem I have but different which I also dont get

OpenStudy (akshay_budhkar):

ok ask here

OpenStudy (anonymous):

I think I can go it just confused on why it has a open dot above the f Given f(x) = 4x + 13 and g(x) = 4x - 1, find (f g)(x).

OpenStudy (akshay_budhkar):

ok now you got to multiply the two functions (f.g)x= f(x).g(x)

OpenStudy (anonymous):

okay same thing just multiplying

OpenStudy (akshay_budhkar):

yea (f+g)x you add f.g you multiply f-g you subtract f/g you divide the functions

OpenStudy (akshay_budhkar):

so u get it now?

OpenStudy (anonymous):

I think so trying to do it now

OpenStudy (akshay_budhkar):

do it if u need any help or get stuck ask for help :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

16x+12?

OpenStudy (akshay_budhkar):

lol i dont know i didnt try

OpenStudy (akshay_budhkar):

dont u know quadratic equations n stuffs?

OpenStudy (anonymous):

Math is my worst subject, good at everything else

OpenStudy (akshay_budhkar):

umm it will take time to teach u everything hehe cant u multiply to linear equations?

OpenStudy (anonymous):

I can do some things lol, I use online calculators to help

OpenStudy (akshay_budhkar):

lol u need to go step by step i dun know how to teach u everything.. i will ask some1 else to as i am leaving now as well lets see if some1 can help you

OpenStudy (anonymous):

thanks its fine, you helped enough :)